A car changes its velocity from 20 m/s [N] to 20 m/s[S] in 5.5 s. If the mass of the car is 1500 kg, 

what is the net force acting on it?



1
Expert's answer
2022-04-11T12:08:44-0400

The Newton's second law says


"F=ma=\frac{m\Delta v}{\Delta t}""F=\frac{1500*(20-(-20))}{5.5}=10909\:\rm N"

The force is directed toward South.